# Structure-preserving and thermodynamically consistent finite element discretization for visco-resistive MHD with thermoelectric effect

### v1

@@ 3.2 Discrete variational formulation
We first consider the case without thermoelectric effect.
@@
$b^{2}_{h}(C,B,u)$ $=\left\langle C,\operatorname{curl}\pi^{\operatorname{curl}}_{h}\left(\pi^{\operatorname{curl}}_{h}B\times\pi^{\operatorname{curl}}_{h}u\right)\right\rangle.$ (46c)
Here $b^{1}_{h}$ is a standard discretization [3] while the choice of $b^{2}_{h}$ is motivated by conservation of magnetic helicity [8].
@@
$g_{h}(w,C,D)$ $=\int_{\Omega}\nu\pi^{\operatorname{curl}}_{h}\operatorname{curl}C\cdot\operatorname{curl}D{\rm d}x.$ (47b)
The operators $d_{h}:F_{h}\times F_{h}\times F_{h}$ and $e_{h}:F_{h}\times F_{h}$ both depend on the prescribed boundary conditions. For homogeneous Neumann conditions, they read
@@
where $\eta>0$ is a penalty parameter. It is a standard non-symmetric interior penalty discretization of the Laplacian [2, § 10.5]. It will be shown, in Section 3.4, to be compatible with the Second Law of Thermodynamics. For non-homogeneous Neumann boundary conditions, they are defined as
@@
The discrete version of (40) is then
@@
In the absence of a magnetic field, the numerical scheme can be derived from a discrete version of the variational principle (19)-(20)-(21) as in [11]. Members of $U^{\operatorname{grad}}_{h}$ are then seen as discrete vector fields that act on functions and densities (i.e. 00-forms and $3$-forms), see [9]. In [4], the action of vector fields is defined on forms of arbitrary order. It would allow to extend the discrete variational principle of [11] to the full system of viscous-resistive MHD. The resulting scheme would only differ from (51) in the expression of $b^{2}_{h}$ so as to reflect the discrete action of vector fields on $1$-forms and $2$-forms.
In practice, it is possible to remove the projection of $\frac{\delta\ell}{\delta u}$ onto $U^{\operatorname{grad}}_{h}$ in (51b) without altering the properties of the scheme. Likewise for the outermost projection of $\pi_{h}\left(w\pi_{h}\frac{\delta\ell}{\delta s}\right)$ in (51e) and the projections $\pi^{\operatorname{div}}_{h}$ inside function $g_{h}$ in (51c) and (51e). Because of the remaining projections, (51) is not directly implementable in this form. In the spirit of [8], intermediate variables are added. For the Lagrangian (30), (51) is then equivalent to finding $u\in U^{\operatorname{grad}}_{h}$, $B\in U^{\operatorname{div}}_{h}$, $\rho,s,\theta,T\in F_{h}$, and $J,H,U,E,\alpha\in U^{\operatorname{curl}}_{h}$ such that
$\langle\partial_{t}(\rho u),v\rangle+a(\rho u,u,v)+b^{1}_{h}(\theta,\rho,v)+b^{1}_{h}(T,s,v)+\langle\alpha,v\rangle=-c(1,u,v),$ $\forall v\in U^{\operatorname{grad}}_{h},$ (52a)
@@
$\langle U,V\rangle=\langle u,V\rangle,$ $\forall V\in U^{\operatorname{curl}}_{h},$ (52i)
$\langle E,F\rangle=-\langle U\times H,F\rangle,$ $\forall F\in U^{\operatorname{curl}}_{h},$ (52j)
$\langle\alpha,\beta\rangle=-\langle J\times H,\beta\rangle,$ $\forall\beta\in U^{\operatorname{curl}}_{h}.$ (52k)
@@ 3.4 Properties of the semi-discrete scheme
$b^{2}_{h}(\pi^{\operatorname{div}}_{h}A,B,u)$ $=\left\langle\pi^{\operatorname{div}}_{h}A,\operatorname{curl}\pi^{\operatorname{curl}}_{h}\left(\pi^{\operatorname{curl}}_{h}B\times\pi^{\operatorname{curl}}_{h}u\right)\right\rangle$ (60a)
$=\left\langle A,\operatorname{curl}\pi^{\operatorname{curl}}_{h}\left(\pi^{\operatorname{curl}}_{h}B\times\pi^{\operatorname{curl}}_{h}u\right)\right\rangle$ (60b)
$=\left\langle\pi^{\operatorname{curl}}_{h}\operatorname{curl}A,\pi^{\operatorname{curl}}_{h}B\times\pi^{\operatorname{curl}}_{h}u\right\rangle$ (60c)
$=\left\langle\pi^{\operatorname{curl}}_{h}B,\pi^{\operatorname{curl}}_{h}B\times\pi^{\operatorname{curl}}_{h}u\right\rangle=0.$ (60d)
@@
Likewise, taking $v=u$ in (52a), $C=-\pi^{\operatorname{div}}_{h}\frac{\delta\ell}{\delta B}$ in (52b), $\sigma=-\pi_{h}\frac{\delta\ell}{\delta\rho}$ in (52c), and $w=1$ in (52d), it eventually gives the total energy rate of change
$\frac{\text{d}}{\text{d}t}\int_{\Omega}\mathcal{E}\,{\rm d}x$ $=\left\langle\partial_{t}\frac{\delta\ell}{\delta u},u\right\rangle-\left\langle\partial_{t}B,\pi^{\operatorname{div}}_{h}\frac{\delta\ell}{\delta B}\right\rangle-\left\langle\partial_{t}\rho,\pi_{h}\frac{\delta\ell}{\delta\rho}\right\rangle-\left\langle\partial_{t}s,\pi_{h}\frac{\delta\ell}{\delta s}\right\rangle,$ (61a)
@@ 3.5 Time discretization
$\hskip 43.36464pt+b^{1}_{h}(T,s_{k+\frac{1}{2}},v)+\langle\alpha,v\rangle=-c(1,u_{k+\frac{1}{2}},v),$ $\forall v\in U^{\operatorname{grad}}_{h},$ (64a)
@@
$\langle U,V\rangle=\langle u_{k+\frac{1}{2}},V\rangle,$ $\forall V\in U^{\operatorname{curl}}_{h},$ (64i)
$\langle E,F\rangle=-\langle U\times H,F\rangle,$ $\forall F\in U^{\operatorname{curl}}_{h},$ (64j)
$\langle\alpha,\beta\rangle=-\langle J\times H,\beta\rangle,$ $\forall\beta\in U^{\operatorname{curl}}_{h}.$ (64k)
@@ 4.1 Reversible flow
This first test case aims at evaluating the performances of the scheme in the absence of dissipation. No viscosity, resistivity or heat exchange are considered (i.e. $\operatorname{Re}=\operatorname{Pm}=\operatorname{Pr}=+\infty$). Gravitation is also removed ($\operatorname{Fr}=+\infty$) and $N$ is set to $0.01$. The heat capacity ratio is chosen as $\gamma=1.4$.
The system under consideration is a square $[0,1]\times[0,1]$ with periodic boundary conditions. The density, temperature and magnetic field are initially homogeneous throughout the entire domain with $\rho=1$, $T=1$ and $B=(0,1)$. A small velocity perturbation $(u_{x},0)$, orthogonal to the magnetic field, is considered and reads
@@
$\frac{\tilde{v}_{c}}{\tilde{v}_{A}}=\sqrt{\frac{\gamma}{N}}\simeq 12.$ (68)
Results for a $20\times 20$ mesh, $r=s=2$ and $\Delta t=0.1$ are given on Figure 1. The evolution of the different energies is consistent with the previous comment. In particular, the ratio of frequencies between the compressive and the magnetic modes is found to be approximately 12. This test case showcases the good behavior of the scheme as oscillations are not visibly damped. This is critical as it means that numerical dissipation is basically absent and will not pollute physical dissipation when it is added.
Figure 1: Evolution of kinetic energy, internal energy and magnetic energy, all integrated over the whole domain. Both internal and magnetic energy evolves approximately as sinusoids and the kinetic energy is the superposition of their opposite. The ratio of frequencies is around 12.
@@ 4.2 Two-dimensional magnetoconvection
The domain $[0,2]\times[0,1]$ is considered with horizontal periodic boundary conditions. At the top and bottom, no-slip boundary conditions are set for the velocity and Dirichlet boundary conditions are enforced on the temperature $T_{\text{bottom}}=1+Z$ and $T_{\text{top}}=1$. Initial conditions follow that of [11] with an additional vertical magnetic field $B=(0,1)$. The temperature is given a linear stratification $T=1+Z(1-z)$. Hydrostatic equilibrium then gives $\rho=T^{m}$ and $p=T^{m+1}$ with $\operatorname{Fr}=\frac{1}{(m+1)Z}$. An initial vertical velocity perturbation $u=(0,u_{z})$ is considered with
@@ 5 Conclusion
The present work paves the way for further studies. The same approach could be used for the numerical approximation of other equations that can be derived from a variational principle. An important example is that of multiphase flows which have a large spectrum of applications in theoretical science and industry alike. Such equations also allow for a virtually infinite number of coupling potentials and dissipation processes, thus making consistency with thermodynamics a critical and challenging target.
@@ Appendix A Consistency of the discrete Seebeck-Peltier coupling
00 0 $=\int_{\Omega}\operatorname{div}(\operatorname{curl}(wT\alpha H))\,{\rm d}x$ (76a)
@@
Furthermore, it is shown in [11, §3.3] that, for smooth $T$, $d_{h}$ satisfies

### v2

@@ 3.2 Discrete variational formulation
The case without thermoelectric effect is first considered.
@@
$b^{2}_{h}(C,B,u)$ $=\left\langle C,\operatorname{curl}\pi^{\operatorname{curl}}_{h}\left(\pi^{\operatorname{curl}}_{h}B\times u\right)\right\rangle.$ (47c)
Here $b^{1}_{h}$ is a standard discretization [4], while the choice of $b^{2}_{h}$ is motivated by conservation of magnetic helicity [10]. Note that in the definition of $b^{1}_{h}$, a centered flux $\{g\}$ was chosen in the boundary term for simplicity. Other choices are also possible; see, for example, [13][§3.5] for a description of how to instead use an upwinded flux without compromising structure preservation.
@@
$g_{h}(w,C,D)$ $=\int_{\Omega}\nu(\pi^{\operatorname{curl}}_{h}\operatorname{curl}C)\cdot\operatorname{curl}D{\rm d}x,$ (48b)
where the curl operator is interpreted in a distributional sense in (48b). The operators $d_{h}:F_{h}\times F_{h}\times F_{h}\rightarrow\mathbb{R}$ and $e_{h}:F_{h}\times F_{h}\rightarrow\mathbb{R}$ both depend on the prescribed boundary conditions. For homogeneous Neumann conditions, they read
@@
where $\eta>0$ is a penalty parameter. It is a standard non-symmetric interior penalty discretization of the Laplacian [3, § 10.5]. It will be shown, in Section 3.4, to be compatible with the Second Law of Thermodynamics. For non-homogeneous Neumann boundary conditions, they are defined as
@@
The discrete version of (41) then consists in finding $u\in U^{\operatorname{grad}}_{h}$, $B\in U^{\operatorname{div}}_{h}$, and $\rho,s\in F_{h}$ such that
@@
In the absence of a magnetic field, the numerical scheme can be derived from a discrete version of the variational principle (19)-(20)-(21) as in [13]. Members of $U^{\operatorname{grad}}_{h}$ are then seen as discrete vector fields that act on functions and densities (i.e. 0-forms and $3$-forms), see [11]. In [5], the action of vector fields is defined on forms of arbitrary order. It would allow to extend the discrete variational principle of [13] to the full system of viscous-resistive MHD. The resulting scheme would only differ from (52) in the expression of $b^{2}_{h}$ so as to reflect the discrete action of vector fields on $1$-forms and $2$-forms.
In practice, it is possible to remove the projection of $\frac{\delta\ell}{\delta u}$ onto $U^{\operatorname{grad}}_{h}$ in (52b) without altering the properties of the scheme. Likewise for the outermost projection of $\pi_{h}\left(w\pi_{h}\frac{\delta\ell}{\delta s}\right)$ in (52e) and the projections $\pi^{\operatorname{div}}_{h}$ inside function $g_{h}$ in (52c) and (52e). Because of the remaining projections, (52) is not directly implementable in this form. In the spirit of [10], intermediate variables are added. For the Lagrangian (31), (52) is then equivalent to finding $u\in U^{\operatorname{grad}}_{h}$, $B\in U^{\operatorname{div}}_{h}$, $\rho,s,\theta,T\in F_{h}$, and $J,H,E\in U^{\operatorname{curl}}_{h}$ such that
$\langle\partial_{t}(\rho u),v\rangle+a(\rho u,u,v)+b^{1}_{h}(\theta,\rho,v)+b^{1}_{h}(T,s,v)-\langle J\times H,v\rangle=-c(1,u,v),$ $\forall v\in U^{\operatorname{grad}}_{h},$ (53a)
@@
$\langle E,F\rangle=-\langle u\times H,F\rangle,$ $\forall F\in U^{\operatorname{curl}}_{h}.$ (53i)
@@ 3.4 Properties of the semi-discrete scheme
$b^{2}_{h}(\pi^{\operatorname{div}}_{h}A,B,u)$ $=\left\langle\pi^{\operatorname{div}}_{h}A,\operatorname{curl}\pi^{\operatorname{curl}}_{h}\left(\pi^{\operatorname{curl}}_{h}B\times u\right)\right\rangle$ (61a)
$=\left\langle A,\operatorname{curl}\pi^{\operatorname{curl}}_{h}\left(\pi^{\operatorname{curl}}_{h}B\times u\right)\right\rangle$ (61b)
$=\left\langle\pi^{\operatorname{curl}}_{h}\operatorname{curl}A,\pi^{\operatorname{curl}}_{h}B\times u\right\rangle$ (61c)
$=\left\langle\pi^{\operatorname{curl}}_{h}B,\pi^{\operatorname{curl}}_{h}B\times u\right\rangle=0.$ (61d)
@@
Likewise, taking $v=u$ in (53a), $C=-\pi^{\operatorname{div}}_{h}\frac{\delta\ell}{\delta B}$ in (53b), $\sigma=-\pi_{h}\frac{\delta\ell}{\delta\rho}$ in (53c), and $w=1$ in (53d), eventually gives the total energy rate of change
$\frac{\text{d}}{\text{d}t}\int_{\Omega}\mathcal{E}\,{\rm d}x$ $=\left\langle\partial_{t}\frac{\delta\ell}{\delta u},u\right\rangle-\left\langle\partial_{t}B,\pi^{\operatorname{div}}_{h}\frac{\delta\ell}{\delta B}\right\rangle-\left\langle\partial_{t}\rho,\pi_{h}\frac{\delta\ell}{\delta\rho}\right\rangle-\left\langle\partial_{t}s,\pi_{h}\frac{\delta\ell}{\delta s}\right\rangle$ (62a)
@@ 3.5 Time discretization
$\hskip 43.36464pt+b^{1}_{h}(T,s_{k+\frac{1}{2}},v)-\langle J\times H,v\rangle=-c(1,u_{k+\frac{1}{2}},v),$ $\forall v\in U^{\operatorname{grad}}_{h},$ (65a)
@@
$\langle E,F\rangle=-\langle u\times H,F\rangle,$ $\forall F\in U^{\operatorname{curl}}_{h}.$ (65i)
@@
Remark 3.4.
Other time discretizations are also possible. Examples include the finite-element-in-time discretization of [1], as well as an extension of the variational time discretization of [11]. The former uses auxiliary variables to systematically preserves invariants and would also allow to retain all properties of the present scheme.
Remark 3.5.
In our implementation, Newton’s method is used to solve system (65) at each time step. Within each iteration of Newton’s method, GMRES is used to solve the linear system. A Schur complement preconditioner is also added, where the velocity is split from all other fields. The inverse of the Schur complement is approximated with a single multigrid V-cycle, and the inverse of the velocity block is approximated with an $ILU(0)$ preconditioner.
Remark 3.6.
Note that equations (65g)-(65i) form a linear systems in the unknowns $J$, $H$, and $E$ which involve non-diagonal mass matrices. To reduce computational expenses, it may be possible to replace these matrices with lumped mass matrices, effectively modifying the inner product on $U_{h}^{\mathrm{curl}}$. However, doing so alters the structure-preserving properties of the scheme; namely, conservation of helicity is lost in the reversible case, while a discrete second principle of thermodynamics is not as immediate in the irreversible case. An interesting topic of future research would be to design mass-lumping schemes that interfere minimally with structure preservation.
@@ 4.1 Reversible flow
This first test case aims at evaluating the performances of the scheme in the absence of dissipation. No viscosity, resistivity or heat exchange are considered (i.e. $\operatorname{Re}=\operatorname{Pm}=\operatorname{Pr}=+\infty$). Gravitation is also removed ($\operatorname{Fr}=+\infty$) and $N$ is set to $0.014$. The heat capacity ratio is chosen as $\gamma=1.4$.
The system under consideration is a square $0,1]\times[0,1]$ with periodic boundary conditions. The density, temperature and magnetic field are initially homogeneous throughout the entire domain with $\rho=1$, $T=1$ and $B=(0,1)$. A small velocity perturbation $(u_{x},0)$, orthogonal to the magnetic field, is considered and reads
@@
$\frac{\tilde{v}_{c}}{\tilde{v}_{A}}=\sqrt{\frac{\gamma}{N}}=10.$ (69)
Results for a $20\times 20$ mesh, $r=s=2$ and $\Delta t=0.1$ are given on Figure 1. The evolution of the different energies is consistent with the previous comment. In particular, the ratio of frequencies between the compressive and the magnetic modes is found to be 10. This test case showcases the good behavior of the scheme as oscillations are not visibly damped. This is critical as it means that numerical dissipation is basically absent and will not pollute physical dissipation when it is added.
Figure 1: Evolution of kinetic energy, internal energy and magnetic energy, all integrated over the whole domain. Both internal and magnetic energy evolves approximately as sinusoids and the kinetic energy is the superposition of their opposite. The ratio of frequencies is 10.
@@ 4.2 Two-dimensional magnetoconvection
The domain $0,2]\times[0,1]$ is considered with horizontal periodic boundary conditions. At the top and bottom, no-slip boundary conditions are set for the velocity and Dirichlet boundary conditions are enforced on the temperature $T_{\text{bottom}}=1+Z$ and $T_{\text{top}}=1$. Initial conditions follow that of [13] with an additional vertical magnetic field $B=(0,1)$. The temperature is given a linear stratification T=1+Z(1z)T=1+Z(1-z). Hydrostatic equilibrium then gives $\rho=T^{m}$ and $p=T^{m+1}$ with $\operatorname{Fr}=\frac{1}{(m+1)Z}$. An initial vertical velocity perturbation $u=(0,u_{z})$ is considered with
@@
Conservation and balance laws.
With the previous Neumann boundary conditions, the top and bottom energy fluxes balance one another so that the total energy is conserved (even though the system is not isolated). The properties of the scheme are then assessed for this test case and results can be found in Figure 5. It is shown that mass and energy are both conserved at the discrete level. So is Gauss’ magnetic law. Finally, the different sources of entropy production are shown to individually have a positive contribution to the total entropy variation.
@@
Figure 5: The properties of the scheme are assessed on the convection test case with Neumann boundary convection and $(\operatorname{Ra},\operatorname{Q})=(4000,10)$. On the second row is plotted the minimum value of the contribution of the different sources of entropy production. Compliance with mass conservation, Gauss’ magnetic law and the two principles of thermodynamics is confirmed at the discrete level.
@@ 5 Conclusion
The present work paves the way for further studies. The same approach could be used for the numerical approximation of other equations that can be derived from a variational principle. An important example is that of multiphase flows which have a large spectrum of applications in theoretical science and industry alike. Such equations also allow for a virtually infinite number of dissipative processes and couplings, thus making consistency with thermodynamics a critical and challenging target. On the other hand, the present method could be adapted to deal with challenging regimes including stiff gradients or general equations of states. Ad hoc numerical modifications would then be needed to ensure positive temperature and densities, as well as accurate and robust shock-capturing.
Acknowledgment.
We are thankful for fruitful discussions with Xavier Garbet. Evan Gawlik was supported by NSF grant DMS-2533499 and the Simons Foundation award MPS-TSM-00002615. François Gay-Balmaz was supported by a startup grant from Nanyang Technological University. François Gay-Balmaz and Bastien Manach-Pérennou were supported by the National Research Foundation Singapore (NRF) core funding ‘Fusion Science for Clean Energy’.
@@ Appendix A Consistency of the discrete Seebeck-Peltier coupling
$0$ $=\int_{\Omega}\operatorname{div}(\operatorname{curl}(wT\alpha H))\,{\rm d}x$ (77a)
@@
Furthermore, it is shown in [13, §3.3] that, for smooth $T$, $d_{h}$ satisfies